OpenStudy (anonymous):

The family functions y=c1*e^x*cos(x)+c2*e^x*sin(x) is a general solution of y''-2y'+2y=0 Decide whether there is a member of the family that satisfies the boundary condition y(-pi/2)=0 and y(pi/2)=0

OpenStudy (turingtest):

we want to look at the possible values of the constants c1 and c2 plug in the values into the equation and you will get a system if the system is solvable for a unique solution of c1 and c2, then that is your answer

OpenStudy (anonymous):

I"ve worked it out, and obviously the cos(x) side of the equation is irrelevant. You end up with two equations 0=c2*e^x*sin(-pi/2) and 0=c2*e^x*cos(-pi/2). Works out to 0= -xe^(-pi/2) and 0=xe^(pi/2)

OpenStudy (anonymous):

I'm having problem expressing it in the form y=(some equation here). I assume there are infinite solutions based a single parameter C. I just can't find what that value is

OpenStudy (turingtest):

\[y(-\frac{\pi}2)=-c_1e^{-\frac\pi2}\cos(-\frac{\pi}2)+c_2e^{-\frac{\pi}2}\sin(-\frac{\pi}2)=-c_2e^{-\frac\pi2}=0\]\[y(-\frac{\pi}2)=-c_1e^{-\frac\pi2}\cos(-\frac{\pi}2)+c_2e^{\frac{\pi}2}\sin(\frac{\pi}2)=c_2e^{\frac\pi2}=0\]so it looks like c2 is trying to be zero, which I'm pretty sure means there is no such equation in the family.

OpenStudy (turingtest):

oh that's all supposed to be positive pi/2 in the second line

OpenStudy (turingtest):

\[y(-\frac{\pi}2)=-c_1e^{-\frac\pi2}\cos(-\frac{\pi}2)+c_2e^{-\frac{\pi}2}\sin(-\frac{\pi}2)=-c_2e^{-\frac\pi2}=0\]\[y(\frac{\pi}2)=-c_1e^{\frac\pi2}\cos(\frac{\pi}2)+c_2e^{\frac{\pi}2}\sin(\frac{\pi}2)=c_2e^{\frac\pi2}=0\]

OpenStudy (turingtest):

hm... you make a point I'm not sure, let me think about it or perhaps @JamesJ can help

OpenStudy (anonymous):

I thought it was 0 at first two, but my online homework disagrees :) thanks for the insight though

OpenStudy (jamesj):

Right. So c1 can be anything and c2 = 0

OpenStudy (turingtest):

Cool, nice and simple thanks James, and that feature is pretty cool :D

OpenStudy (jamesj):

Yes, it is. I just hope no every man and his dog can mention me!

OpenStudy (anonymous):

Also good, my question is how would I write that in terms of y=______ , though I agree your answer is much more intuitive

OpenStudy (jamesj):

*not

OpenStudy (turingtest):

I hear you brycoop: I believe you can write if you want to express that you can write\[y=te^x\cos x;t\in \mathbb R\]

OpenStudy (jamesj):

y=c1*e^x*cos(x)

OpenStudy (turingtest):

or yeah, same thing...

OpenStudy (anonymous):

ehh webworks won't take it, which doesn't mean its wrong, I'll email my prof and see what she says, thanks a bunch guys

OpenStudy (turingtest):

perhaps your webwork would be happier if you chose a random number for c, like c=1 have you tried that?

OpenStudy (anonymous):

just needed to add a * forget how picky webworks can be with multiplication. Try to avoid being implicit in the future I suppose haha